### v2

# Toward Universal Laws of Outlier Propagation

Abstract
We argue that Algorithmic Information Theory (AIT) admits a principled way to quantify outliers in terms of so-called randomness deficiency. For the probability distribution generated by a causal Bayesian network, we show that the randomness deficiency of the joint state decomposes into randomness deficiencies of each causal mechanism, subject to the Independence of Mechanisms Principle. Accordingly, anomalous joint observations can be quantitatively attributed to their root causes, i.e., the mechanisms that behaved anomalously. As an extension of Levin’s law of randomness conservation, we show that weak outliers cannot cause strong ones when Independence of Mechanisms holds. We show how these information theoretic laws provide a better understanding of the behaviour of outliers defined with respect to existing scores.
**footnotetext: These authors contributed equally to this work.
1 Introduction
Anomaly detection plays a crucial role in business, technology, and medicine. Typical use cases range from fraud detection in finance and online trading [Donoho, 2004], performance drops in manufacturing lines [Susto et al., 2017] or cloud computing applications [Gan et al., 2021, Ma et al., 2020, Hardt et al., 2023], health monitoring in intensive care units [Maslove et al., 2016], to explaining extreme weather and climate events [Zscheischler et al., 2022]. It has motivated a vast effort towards developing methodologies relevant to outlier analysis. As example, we refer the reader to some of the early works in statistics and computer science [Freeman, 1995, Rocke and Woodruff, 1996, Rousseeuw and Leroy, 2003, Aggarwal, 2017]. In complex systems, an anomaly will typically cause a large cascades of related anomalies [Panjei et al., 2022]. In order to mitigate them, it is not sufficient to merely detect the anomalies; we must also identify which of the anomalies was the root cause [Budhathoki et al., 2022, Ikram et al., 2022, Li et al., 2022, Hardt et al., 2023, Wang et al., 2023b, a]. Thus, we implicitly face the counterfactual question of what conditions could have been different to prevent the (usually undesired) anomalous event.
To render a complex system accessible to human understanding, we begin with a causal model of its relevant mechanisms, specifying not only their default behavior, but also their behavior under modifications called interventions. Such a model should be modular in two respects. First, we may want to understand the causal pathway, along which a perturbation of any part of the system propagates through its components until it generates the event. Second, we want to “blame” some component(s) of the system, while acknowledging that others worked as expected.
Causal Bayesian networks offer a framework that supports both kinds of modular description, specifying causal relations via a directed acyclic graph (DAG) $G$ with random variables $X_{1},\dots,X_{n}$ as nodes [Pearl, 2009, Spirtes et al., 1993]. Due to the causal Markov condition [Pearl, 2009], the joint distribution factorizes according to
$P(X_{1},\dots,X_{n})=\prod_{i=1}^{n}P(X_{j}\mid\operatorname{{PA}}_{j}),$ (1)
where $\operatorname{{PA}}_{j}$ denotes the parents of $X_{j}$ in $G$, i.e., its direct causes. We will think of each conditional distribution $P(X_{j}\mid\operatorname{{PA}}_{j})$ as an independent mechanism of the system, which can in principle be changed or replaced without changing the others (see 2.1 and 2.2 in Peters et al. [2017] for a historic overview). Following this modular view, our root cause analysis will assume that in the case of an anomalous observation, most of the mechanisms worked as expected; thus, the anomaly can be blamed on a small number of mechanisms that act as root causes, in alignment with Schölkopf et al. [2021]’s “sparse mechanism shift hypothesis”.
To our knowledge, Budhathoki et al. [2022] provide the most elaborate formalization of the idea of attributing anomalies to mechanisms. We develop our concepts starting from this baseline.
1.1 Outlier scores from p-values
To quantitatively attribute an anomalous event to upstream nodes, Budhathoki et al. [2022] first introduce what they call an Information Theoretic (IT) outlier score via
$\lambda_{\tau}(x):=-\log P(\tau(X)\geq\tau(x)),$ (2)
where $x$ denotes an observation of the random variable $X$, and $\tau:{\cal X}\to\mathbb{R}$ is an appropriate feature statistic, whose choice will be discussed later [^1] While Budhathoki et al. [2022] use the natural base $e$, we use base $2$ logarithms to align with binary program lengths in algorithmic information theory. In effect, we express the outlier score in units of bits [Frank, 2005]..
$\lambda_{\tau}$ can be viewed as a statistical test of the null hypothesis that $x$ was sampled from $P$: setting the base of the logarithm to $2$ yields the p-value $2^{-\lambda_{\tau}(x)}$. A small p-value (or large $\lambda_{\tau}$) corresponds to an unusual sample under $P$, which can be labeled an outlier. Since $x$ is a single observation, anomaly scoring thus reduces to classical hypothesis testing on a sample size of $1$ [Shen, 2020, Vovk, 2020] [^2] Note that Tartakovsky et al. [2012] discuss anomalies as change points over multiple observations. However, we focus on anomalies confined to an individual observation..
1.2 Quantitative root cause analysis in causal Bayesian networks
To quantitatively attribute an anomalous observation $(x_{1},\dots,x_{n})$ to different mechanisms, where each $x_{j}$ represents a value of the corresponding variable $X_{j}$, the arXiv version of Budhathoki et al. [2022] define conditional outlier scores:
$\lambda_{\tau}(x_{j}\mid\operatorname{{pa}}_{j}):=-\log P(\tau(X_{j})\geq\tau(x_{j})\mid\operatorname{{PA}}_{j}=\operatorname{{pa}}_{j}).$
They demonstrate that this score can be equivalently interpreted as measuring the outlierness of the noise term [^3] This aligns in spirit with Von Kügelgen et al. [2023], where events are also backtracked by attributing them to the noise variables.. The feature functions $\tau_{j}$ can be node-specific, which is essential when the variables $X_{j}$ have different characteristics (e.g., different dimensionality or data types). When the specific choice of $\tau$ is not crucial for the discussion, we will simplify notation by dropping the subscript. The arXiv version of Budhathoki et al. [2022] extends this framework by introducing a joint outlier score through “convolution”:
$\lambda(x_{1},\dots,x_{n})$ ::=:= $\sum_{j=1}^{n}\lambda(x_{j}\mid\operatorname{{pa}}_{j})$
-- $\log\sum_{i=1}^{n-1}\frac{(\sum_{j=1}^{n}\lambda(x_{j}\mid\operatorname{{pa}}_{j}))^{i}}{i!}.\vspace{-0.2cm}$
The second term serves as a correction factor that ensures the score maintains the properties of an IT score, provided that the conditional distributions have densities with respect to the Lebesgue measure.
1.3 Monotonicity of scores
Okati et al. [2024] formalize an intuitive principle about anomaly propagation: unless the connecting mechanisms themselves are also anomalous, a moderate outlier in a cause (measured by $\lambda(x_{1})$) should not produce an extreme outlier in its effect (measured by $\lambda(x_{2})$). This can be demonstrated for the bivariate causal relation $X_{1}\to X_{2}$ as follows:
Lemma 1.1 (Weak IT outliers rarely cause strong ones).
If $\lambda$ are IT scores and $X_{1}$ is a continuous variable, the following inequality holds:
$P(\lambda(X_{2})\geq\lambda(x_{2})\mid\lambda(X_{1})\geq c)\leq 2^{c-\lambda(x_{2})}.$
In other words, whenever we generate an anomaly $x_{1}$ randomly from $P(X_{1}\mid\lambda(X_{1})\geq c)$ and then generate $X_{2}$ via its mechanism $P(X_{2}\mid X_{1})$, the resulting outlier score is unlikely to be much larger than $c$. The intuitive reason is that the probability occupied by events with high scores is much smaller than that of a given event having a lower score.
Note that there is no statement for any single $x_{1}$ – some of them may generate outliers that are higher with large probability – but “most” of them do not. The reason that we cannot make any statement about particular values is crucial for the present paper, because it shows how independence of mechanisms comes into play: whenever one chooses, on purpose, an $x_{1}$ such that it results in a large outlier downstream, there is no reason that this particular $x_{1}$ needs to have high outlier score. One can easily think, for instance, of the deterministic relation $X_{2}=f(X_{1})$, where $f$ is a highly non-linear function that results in rather anomalous values $x_{2}$ for some tiny region of values $x_{1}$ which are not anomalous.
1.4 Limitations of current approach
Rigid definition of outliers
$\lambda_{\tau}$ defines outliers in terms of a feature statistic $\tau$ that is chosen in advance. Sometimes, we only know what makes a sample abnormal after seeing it, so we would like to be able to choose $\tau$ with hindsight. For example, suppose $P$ is univariate Gaussian. It seems natural to define $\tau(x):=|x-\mu|$, so that the anomaly detector flags extremely high or low values. However, we might also like to flag observations such as $x=0$ or $x=\mu$: while such $x$ may lie in a region of high density, the probability of obtaining such a specific value is zero under $P$. Although any other value has also probability zero, one may agree that observing $x=0$ and $x=\mu$ would be surprising because these values are special (in a sense specified later).
As another example, suppose each component of a multivariate observation $x$ is the reading of a different sensor. The signal transmission from all $d$ sensors may be broken in such a way that all components equal to the same constant “idle” state $c$. Such coincidences also indicate an unusual event.
Aggarwal [2016] describe a broad variety of different outliers beyond the above toy examples: These can be unusual frequencies of words in text documents [Mohotti and Nayak, 2020], unexpected patterns in images [Yakovlev et al., 2021], unusually large cliques in graphs [Hooi et al., 2016], or points lying in low density regions [Breunig et al., 2000].
No general decomposition rule
While (1.2) nicely decomposes the joint outlier score into mechanism-specific conditional scores, a key limitation is that this decomposition relies on defining the joint score based on a sum of conditional scores. This does not imply that any reasonable outlier score (e.g., using a generic feature function $\tau$) for the joint observation can be decomposed in this manner.
1.5 Our contributions
The contribution of this paper is purely conceptual. We do not propose another outlier detection or root cause analysis method, but instead provide a theoretical framework for calibrating and interpreting outlier scores. The framework ensures that outlier scores meet three critical criteria: (i) Comparability across diverse probability spaces and data modalities. (ii) Non-increasingness along causal chains of downstream effects, regardless of variable modalities. (iii) Well-defined attribution of joint system outlier scores to anomalies of mechanisms.
Our ideas were guided by the following general working hypothesis, which we believe applies far beyond the subject of this paper:
Principle 1.2 (Information Theory as a Guide).
Good information theoretic concepts enable many nice theorems, but they are often hard to work with in practice. However, together with distributional assumptions (e.g. Gaussianity) they can boil down to simple concepts (e.g. linear algebraic expressions). The resulting formulae may be valid and useful beyond the distributional assumptions (e.g. by virtue of linear algebra).
While Shannon information is sometimes hard to estimate from small sample sizes, algorithmic information is even worse: Kolmogorov complexity is not even computable. Furthermore, its identities hold only up to machine-dependent additive constants. Therefore, we owe it to the reader to show that our algorithmic information theoretic concepts trigger insights that can be applied in practice, as we will try in Section 5.
The paper is structured as follows. Section 2 describes concepts from statistics, information theory, and causality that we build upon. Section 3 derives the decomposition, and Section 4 shows that the joint score is non-increasing under marginalization. Section 5 discusses simple examples and Section 6 a toy experiment. For a cleaner exposition, we defer some formal proofs and definitions to the supplementary material.
2 Key ingredients
2.1 Statistical testing with e-values instead of p-values
While p-values are the most famous measure of evidence in statistical testing, e-values are recently gaining popularity for their superior ability to aggregate evidence across multiple tests [Ramdas and Wang, 2024]. These values are inconsistently scaled in the literature, with e-values being comparable to reciprocals of p-values and exponentials of algorithmic randomness scores. To remove any obfuscation coming from scaling conventions, we introduce the following definitions.
Definition 2.1.
A probability-bounded test (p-test) in ratio form is a statistic $\Lambda:\mathcal{X}\rightarrow[0,\infty]$ satisfying $\forall\varepsilon>0,P\left(\Lambda(X)\geq 1/\varepsilon\right)\leq\varepsilon.$ An expectation-bounded test (e-test) in ratio form is a statistic $\Lambda:\mathcal{X}\rightarrow[0,\infty]$ satisfying $\operatorname*{\mathbf{E}}_{X\sim P}\left(\Lambda(X)\right)\leq 1.$
We say a statistic $\Lambda:\mathcal{X}\rightarrow[0,\infty]$ is a p-test (or e-test) in probability form, if $1/\Lambda$ is a p-test (or e-test) in ratio form. Note that what is commonly called a “p-value” is a p-test in probability form, satisfying $\Lambda(X)\leq\epsilon$ with probability at most $\epsilon$. In contrast, what the literature calls an “e-value” is an e-test in ratio form.
Similarly, a statistic $\lambda:\mathcal{X}\rightarrow[-\infty,\infty]$ is a p-test (or e-test) in log form, if $2^{\lambda}$ is a p-test (or e-test) in ratio form. Our convention is to use lowercase symbols like $\lambda$ to indicate log form.
By Markov’s inequality, every e-test is also a p-test. Conversely, Vovk and Wang [2021] describe a number of ways to calibrate any given p-test into an e-test. For more details on tests, see Section 8 in the supplementary material.
p-tests provide a straightforward way to control the Type I error rate, i.e., false positives. To ensure that samples from $P$ are labeled as anomalies at a rate no larger than a desired threshold $\epsilon$, we flag only those samples whose p-test scores are above $1/\epsilon$ (i.e., below $\epsilon$ when expressed in probability form).
Since every e-test is also a p-test, e-tests also achieve this false positive rate; however, they are more conservative. In return, e-tests offer many advantages related to composability and optional stopping [Grünwald et al., 2020, Ramdas et al., 2023, Ramdas and Wang, 2024], and we will find them more convenient for decomposing anomaly scores by mechanism.
2.2 Basic notions from algorithmic information theory
The intuition behind universal tests is that an observation is anomalous precisely when it is more compressible than usual for the underlying random process. To formalize description lengths, we fix a universal prefix-free Turing machine. The conditional Kolmogorov complexity $K(x\mid y)$ is the bit length of the shortest program $p$ that outputs $x$ when given access to another string $y$ [Li and Vitányi, 2019]. It satisfies the Kraft inequality $\sum_{x}2^{-K(x\mid y)}<1$. When $y$ is an empty string, we write $K(x)$. Just as Shannon’s entropy measures information content for a probability distribution, $K(x)$ measures it for an individual sample $x$.
Using a standard prefix-free encoding of $n$-tuples, Li and Vitányi [2019] also define the joint Kolmogorov complexity $K(x_{1},\dots,x_{n})$. By analogy with Shannon’s mutual information, they define the algorithmic mutual information between $x$ and $y$, conditional on $z^{*}$, by
$I(x:y\mid z^{*})$ $:=K(x\mid z^{*})+K(y\mid z^{*})-K(x,y\mid z^{*})$
=$\stackrel{{\scriptstyle+}}{{=}}K(x\mid z^{*})-K(x\mid(y,z)^{*}).$
Here, $z^{*}$ denote a shortest program that outputs $z$, and $\stackrel{{\scriptstyle+}}{{=}}$ denotes equality up to a constant dependent on the universal machine, but not on $x$ or $y$ [^4] Note that $x^{*}$ contains more information than $x$, because $x$ is easily generated from $x^{*}$ but not vice versa. Conditioning on $x$ instead of $x^{*}$ would result in error terms that are often constant, and at worst logarithmic in the length of $x$ [Gács, 2021].. We say $x$ and $y$ are conditionally independent, given $z$, if $I(x:y\mid z^{*})\stackrel{{\scriptstyle+}}{{=}}0$.
2.3 Universal tests
A Martin-Löf or Levin test (i.e., semicomputable p-test or e-test, respectively) can be thought of as combining all computable anomaly scoring algorithms, in order to detect the broadest possible variety of anomalies. Section 9 in the supplementary material contains full details; here we give a brief overview.
Definition 2.2 (Domination property).
For two statistical tests $\lambda_{1}$ and $\lambda_{2}$ expressed in log form, we say $\lambda_{1}$ dominates $\lambda_{2}$ if there exists a constant $c\in\mathbb{R}$ such that for all observations $x$ in the sample space,
$\lambda_{1}(x)\geq\lambda_{2}(x)-c.$
Intuitively, this means $\lambda_{1}$ can detect at least all the anomalies that $\lambda_{2}$ can detect, up to a constant term.
The domination property provides a natural way to compare the power of different statistical tests. The class of semicomputable tests contains a universal test that dominates all the others:
Theorem 2.3 (Universality of randomness deficiency).
Let $\mathcal{X}$ be a discrete space that can be interpreted as a subset of $\{0,1\}^{*}$ in a canonical way, and $P$ be a computable probability distribution on $\mathcal{X}$ (i.e., with finite description length). Then, the randomness deficiency of $x\in\mathcal{X}$, defined by
$\delta(x):=-\log P(x)-K(x\mid P^{*}),$ (4)
is a universal e-test, dominating all other semicomputable e-tests.
The intuition behind Eq. 4 is that typical samples from a distribution $P$ are optimally compressed by encodings of length $-\log P(x)$. When a sample $x$ can be compressed beyond this theoretical limit, we consider it “anomalous”. This includes the case where we observe $x=0$ from a discretized centered Gaussian distribution. Despite being the mode of the distribution, 00 is considered anomalous because its negative log-likelihood $-\log P(x)$ is substantially larger than its Kolmogorov complexity $K(0)\stackrel{{\scriptstyle+}}{{=}}0$.
Note that when a distributional estimate $\hat{P}$ is inferred from data, the “reconstruction loss” $-\log\hat{P}(x)$ can be used as an outlier score [Bishop, 1993]. Without further considerations, this score is not calibrated because there may be a huge region with low density, and thus all low density points together may still be likely. However, if the term $K(x\mid P^{*})$ is also small, we know that $x$ is special within this huge set. [^5] Note that conformal p-values [Bates et al., 2021] do not require any parametric assumptions or density estimation, since their coverage guarantees rely on exchangeability alone. However, statements on conditional coverage [Barber et al., 2019] are too weak for our purpose of calibrating conditional outlier scores.
Example 1 (Uniform distribution).
When $P$ is uniform over all $2^{d}$ binary strings of length $d$, 4 becomes $\delta(x)=d-K(x\mid P^{*})\stackrel{{\scriptstyle+}}{{=}}d-K(x\mid d)$, measuring the degree to which $x$ is compressible.
2.4 Algorithmic Markov condition and independence of mechanisms
Our goal is to demonstrate how the randomness deficiency decomposes across causal mechanisms in a Bayesian network, subject to postulates that connect algorithmic information theory to causality. To achieve this, we establish a theoretical framework based on fundamental postulates.
Janzing and Schölkopf [2010] propose an adaptation of (1) that characterizes algorithmic dependencies between individual objects, rather than statistical dependencies between random variables:
Postulate 2.4 (Algorithmic Markov condition).
Let $x_{1},\dots,x_{n}$ be binary words describing objects whose causal relations are given by the DAG $G$. Then the joint complexity of $(x_{1},\dots,x_{n})$ decomposes as
$K(x_{1},\dots,x_{n})\stackrel{{\scriptstyle+}}{{=}}\sum_{j=1}^{n}K(x_{j}\mid\operatorname{{pa}}^{*}_{j}).$
Furthermore, for any three sets $R,S,T$ of nodes, we have
$I(R:S\mid T^{*})\stackrel{{\scriptstyle+}}{{=}}0,$
whenever $R$ and $S$ are d-separated by $T$.
Janzing and Schölkopf [2010] argue that within a causal Bayesian network, where nodes represent random variables, the mechanisms $P_{X_{j}|\operatorname{{PA}}_{j}}$ constitute independent information-bearing objects. Given $m$ observations of $n$-tuples $(x^{1}_{1},\dots,x_{n}^{1}),\dots,(x^{m}_{1},\dots,x_{n}^{m})$, they define a DAG $G^{m}$ that connects the $n$ mechanisms $P_{X_{i}|\operatorname{{PA}}_{i}}$ with the $n\times m$ observations in the following way: Each $x_{i}^{l}$ has the observations $\operatorname{{pa}}^{l}_{i}$ and the node $P_{X_{i}|\operatorname{{PA}}_{i}}$ as parents, formalizing the idea that the mechanisms $P_{X_{i}|\operatorname{{PA}}_{i}}$ also determine how the parents in $G$ influence the respective observation. Figure 1 shows $G^{m}$ for the DAG $X\to Y$. For our purpose, it is sufficient to consider $G^{1}$, where we have only one observation $x_{j}$ from each node $X_{j}$ in $G$. Accordingly, we construct $G^{1}$ from $G$ as follows: replace each variable $X_{j}$ with its observation $x_{j}$, and make $P_{X_{j}|\operatorname{{PA}}_{j}}$ a parent of $X_{j}$. Each $P_{X_{j}|\operatorname{{PA}}_{j}}$ becomes a root node, and these are the only root nodes in $G^{1}$. This placement of $P_{X_{j}|\operatorname{{PA}}_{j}}$ as root nodes (and thus algorithmically independent) formalizes the Independence of Mechanisms [^6] Janzing and Schölkopf [2010] call the algorithmic independence of different $P_{X_{j}\mid\operatorname{{PA}}_{j}}$ Independence of Mechanisms. Here we prefer using this term for the underlying idea of representing them as root nodes.. Applying the algorithmic Markov condition to $G^{1}$ yields:
Lemma 2.5 (Conditional irrelevance of other mechanisms and predecessors when parents are given).
Let $X_{1},\dots,X_{n}$ be causally ordered. Given its parents $\operatorname{{pa}}_{j}$ and the mechanism $P_{X_{j}\mid\operatorname{{PA}}_{j}}$, none of the other mechanisms $(P_{X_{i}\mid\operatorname{{PA}}_{i}})_{i\neq j}$ and none of the causal predecessors $(x_{i})_{i<j}$ enable further compression of $x_{j}$. That is,
[…]
3 Decomposition of Randomness Deficiency
We start by presentation the decomposition of randomness deficiency in the bivariate case. Specifically, consider the DAG $X\rightarrow Y$ between a cause $X$ and its effect $Y$. Extending Eq. 4 with each distribution’s shortest representation, define the joint randomness deficiency of outcomes $(x,y)$ with respect to $P_{X,Y}$ by
$\delta(x,y)\vcentcolon=-\log P(x,y)-K(x,y\mid(P_{X,Y})^{*}),$ (5)
and the conditional randomness deficiency by
$\delta(y\mid x)\vcentcolon=-\log P(y\mid x)-K(y\mid(x,P_{Y\mid X})^{*}).$ (6)
Lemma 3.1.
(Decomposition of randomness deficiency for a cause-effect pair) For any two random variables $X\rightarrow Y$ (i.e., $X$ being the cause of $Y$), and for specific observations $x$ and $y$, the following equality holds under Postulate 2.4 for the DAG in Figure 1:
$\delta(x,y)\stackrel{{\scriptstyle+}}{{=}}\delta(x)+\delta(y\mid x)$
The proof relies on Lemma 2.5, and is provided in Section 10 of the supplementary.
[…]
Figure 1: $x_{1},\dots,x_{m}$ sampled from $P_{X}$, $y_{1},\dots,y_{m}$ sampled from $P_{Y\mid X}$.
The following example shows why causal assumptions are needed for randomness deficiency to be additive.
Example 2 (No additivity without Independence of Mechanisms).
Let $\mathcal{X}=\mathcal{Y}=\{0,1\}^{d}$ and $P_{X}$ be the uniform distribution. Furthermore, let $P_{Y|X}$ be the deterministic mechanism
$P(y\mid x)=1\quad\hbox{ iff }\quad y=x\oplus x_{0},$
where $\oplus$ denotes bitwise XOR and $x_{0}$ is an algorithmically random string with $K(x_{0})=d$. The mechanism $P_{Y\mid X}$ being deterministic implies that we always have $\delta(y\mid x)\stackrel{{\scriptstyle+}}{{=}}0$.
[…]
We now generalize Lemma 3.1 to the multivariate case:
Theorem 3.2 (Decomposition of multivariate joint randomness deficiency).
Let the set of strings $x_{1},x_{2},\dots,x_{n}$ be causally connected by a directed acyclic graph $G$, so that the causal Markov condition holds for $G^{m}$. Then the joint randomness deficiency of all strings $x_{1},x_{2},\ldots,x_{n}$ decomposes into the conditional randomness deficiencies of the mechanisms:
$\delta(x_{1},\dots,x_{n})\stackrel{{\scriptstyle+}}{{=}}\sum_{j=1}^{n}\delta(x_{j}\mid\operatorname{{pa}}_{j}),$
where $\delta(x_{j}\mid\operatorname{{pa}}_{j})$ denotes the randomness deficiency of $x_{j}$ given its parents.
The proof is a simple induction over the number $n$ of nodes according to the causal ordering. The induction step from $n-1$ to $n$ uses Lemma 3.1 where $y:=x_{n}$ and $x$ is the single multivariate parent $x:=(x_{1},\dots,x_{n-1})$; see Section 10 in the supplementary for details.
4 Monotonicity of Randomness Deficiency
In particular, Theorem 3.2 allows us to draw the following straightforward conclusion:
Theorem 4.1 (Weak anomalies do not cause stronger ones).
If there is a unique root cause $j\in\{1,\dots,n\}$, in the sense that
$\delta(x_{i}\mid\operatorname{{pa}}_{i})\stackrel{{\scriptstyle+}}{{=}}0\quad\hbox{ for }i\neq j,$
and the conditions of Theorem 3.2 are met,
[…]
The proof follows easily from Theorem 3.2 together with Corollary 4.1.11 from Gács [2021] which states non-increasingness of $\delta$ under marginalization and thus
$\delta(x_{i})\stackrel{{\scriptstyle+}}{{\leq}}\delta(x_{1},\dots,x_{n}).$
In essence, (7) states that none of the nodes displays a randomness deficiency that exceeds the conditional randomness deficiency of the root cause. In a nutshell, “weak outliers cannot cause strong ones”. – We have thus found an AIT version of Lemma 1.1.777Hence, in the language of modern resource theories in physics [Coecke et al., 2016], Definition 5.1., an anomaly is a “resource” and $\delta$ a “monotone” measuring its value. It is worth noting that Levin [1984] (see also Gács [2021]) demonstrated a similar principle, calling it randomness conservation. It says that the output of a mechanism cannot exhibit a substantially larger randomness deficiency than the input - under the condition that the mechanism itself has a constant description length. In physics, randomness deficiency corresponds to a lack of entropy [Zurek, 1989, Gács, 1994]. Therefore, the second law of thermodynamics amounts to an instance of randomness conservation [Ebtekar and Hutter, 2025].
In our context, simplicity of mechanisms is generalized by the independence of mechanisms. Example 2 illustrates why the independence of mechanisms principle is essential. It describes a scenario where a value $y$ with randomness deficiency $d$ emerges, yet this deficiency cannot be attributed to either the input $x$ or the mechanism generating $y$ from $x$.
5 Relation to computable anomaly scores
This section describes simple scenarios in which randomness deficiency boils down to simple and well-known scores. One may ask why one should start with an uncomputable concept in the first place to end up with something simple anyway. The answer is that the properties we have shown for randomness deficiency guide us in calibrating simple scores, with desirable properties such as decomposition of joint scores into scores of the mechanisms and monotonicity of scores.
Example 3 (z-score).
For a Gaussian variable $X\sim{\mathcal{N}}(\mu,\sigma)$, the squared z-score reads $z^{2}(x):=(x-\mu)^{2}/\sigma^{2}.$ At a fixed level of precision, $K(x\mid P^{*})\stackrel{{\scriptstyle+}}{{\leq}}2\log|x-\mu|$. Substituting the log likelihood
$-\log P(x)$ $=\log\sqrt{2\pi}\sigma+\frac{\log e}{2}z^{2}(x),$
and treating $\sigma$ as a constant, yields
$\delta(x)$ $\stackrel{{\scriptstyle+}}{{\geq}}\frac{\log e}{2}z^{2}(x)-2\log|x-\mu|.$
Note that while the well-known identity $\mathbf{E}[z^{2}]=1$ tells us that $z^{2}$ is an e-value, the universal e-value $2^{\delta}\approx e^{z^{2}/2}$ provides far stronger evidence at the tails. Now we generalize to $n>1$ dimensions:
Example 4 (Decomposing squared Mahalanobis distance).
The randomness deficiency of a random vector $\textbf{x}\in\mathbb{R}^{n}$, drawn from a centered Gaussian, satisfies
$\delta(\textbf{x})\stackrel{{\scriptstyle+}}{{\geq}}\frac{\log e}{2}\textbf{x}^{T}\Sigma^{-1}_{\textbf{X}}\textbf{x}-O(\log\|\textbf{x}\|_{\infty}),$
where $\Sigma_{\textbf{X}}$ denotes the covariance matrix and $n$ is treated as a constant. The leading term is $(\log e)/2$ times the squared Mahalanobis distance (M-distance, for short).
Let X be generated by a causal Bayesian network with linear structural equations $\textbf{X}=A\textbf{X}+{\bf N}$, where $A$ is strictly lower triangular and $N_{i}$ are independent noise variables with variance $\sigma_{i}^{2}$. Using transformation ${\bf N}=(I-A)\textbf{X}$, we obtain a diagonal form
$\textbf{x}^{T}\Sigma^{-1}_{\textbf{X}}\textbf{x}=\sum_{i=1}^{n}\frac{n_{i}^{2}}{\sigma_{i}^{2}}.$
For large $n_{i}$, the expression $n_{i}^{2}/\sigma_{2}$ is roughly proportional to the randomness deficiency of the mechanism $P(X_{i}\mid PA_{i})$. Hence, decomposition of randomness deficiencies translates asymptotically into decompositions of the squared M-distance (as used for multivariate anomaly detection Aggarwal [2016]) into $z^{2}$-scores. Further, M-distance is non-increasing with respect to marginalization to a subset of variables (see Section 11 in the supplementary), resembling the monotonicity of randomness deficiency.
Example 4 supports Principle 1.2: a conservative bound on the randomness deficiency provides us with a computable e-value, as a function of the $z^{2}$-scores. Moreover, the conclusion that the $z^{2}$-score of any variable cannot exceed the sum of all these “noise scores” still holds, because our reasoning above is entirely based on linear algebra.
The following example shows that root causes may have smaller marginal randomness deficiency than their downstream effect:
Example 5 (Root cause with small marginal score).
Let $X_{2}=2X_{1}+E_{2}$ with $X_{1},E_{2}\sim{\mathcal{N}}(0,1)$. Then $X_{2}\sim{\mathcal{N}}(0,\sqrt{3})$. Increasing $x_{2}$ by manual perturbation induces a marginal randomness deficiency that increases with $x_{2}^{2}/3$ for $x_{2}\to\infty$, while the conditional deficiency increases even like $x^{2}_{2}/1$. Thus, we have $\delta(x_{2}\mid x_{1})>\delta(x_{2})$ and $\delta(x_{1},x_{2})>\delta(x_{2})$ for large perturbations. If $X_{1}$ and $X_{2}$ together influence a third variable $X_{3}=X_{1}-X_{2}+E_{3}$ with $E_{3}\sim\mathcal{N}(0,1)$, we have $X_{3}\sim{\mathcal{N}}(0,\sqrt{2})$, thus $\delta(x_{3})$ increases with $x^{2}_{2}/2$. Hence, $\delta(x_{3})$ increases stronger than $\delta(x_{2})$, but weaker than $\delta(x_{2}\mid x_{1})$ and $\delta(x_{1},x_{2})$.
This paradox resembles Example 2.1 in Li et al. [2024], where it is phrased in terms of $z^{2}$-scores. Following Principle 1.2, we can describe it entirely in terms of M-distances (which are just $z^{2}$-scores in the one-dimensional case): the M-distance of $x_{3}$ can be larger than the M-distance of its root cause $x_{2}$, but not larger than the M-distance of the pair $(x_{1},x_{2})$, that is, the full cause of $x_{3}$. In other words, to fully quantify the anomaly introduced by perturbing $x_{2}$ we need to also account for the destroyed coupling between $x_{1}$ and $x_{2}$, not only for the size of the value $x_{2}$ itself.
The merit of algorithmic information theory is the insight that anomaly scores of the vector $(x_{1},x_{2})$ and the scalar $x_{3}$ are indeed comparable, and also comparable to the conditional score of $x_{2}$ given $x_{1}$, because their calibration was guided by the randomness deficiency.
In the following example, which could be easily generalized to words over an arbitrary alphabet, randomness deficiency essentially boils down to relative entropy:
Example 6 ($m$-bit binary word).
Let us first consider an $m$-bit word whose bits are set to $1$ independently with probability $p$. Counting the number of words $w$ with fixed Hamming weight $\ell$ yields the lower bound
$\delta(w)\stackrel{{\scriptstyle+}}{{\geq}}-\ell\log p-(m-\ell)\log(1-p)-\log{m\choose\ell}-O(\log m).$
Using Stirling’s approximation $\log m!=m\log m-m\log e+O(\log m)$, one can show that
$\delta(w)\stackrel{{\scriptstyle+}}{{\geq}}m\cdot{\operatorname{KL}\left({\frac{\ell}{m}\mid\mid p}\right)}-O(\log m),$ (8)
with the binary Kullback-Leibler distance
${\operatorname{KL}\left({q\mid\mid p}\right)}:=q\log\frac{q}{p}+(1-q)\log\frac{1-q}{1-p}.$
Here we can see that words with unexpectedly low or high Hamming weight, both are assigned high outlier scores, enabling us to compare both types of outliers. This works despite the probability mass function being monotonic in the Hamming weight $\ell$.
Multiple works [Aggarwal, 2016, Akoglu et al., 2012, Noble and Cook, 2003] propose detecting anomalies by compression, but consider anomalies that have higher compression length than usual, although randomness deficiency seems to flag only anomalies that have unexpectedly low compression length. Example 6 also resolves this paradox: For $p<1/2$, any string with $K(w)>m\cdot H_{2}(p)$ must have Hamming weight larger than $mp$. Accordingly, its randomness deficiency is bounded from below by (8). Hence, paradoxically, an unusually high compression length also implies non-zero randomness deficiency, because it entails an increase of the log likelihood term that outweighs the observed increase in compression length.
More sophisticated notions of anomalies can be obtained, for instance, when $X$ is graph-valued and an anomaly is given by large cliques [Aggarwal, 2016] – e.g. when fraudsters work together frequently on illegal activities, their communication is densely connected. To estimate the random deficiency of a graph $G$ with a clique of size $k$, one only needs to define a probability distribution on the set of graphs with $m$ nodes, and bound the description length of a graph by counting the number of graphs with cliques of size $k$.
6 Experiment with Lempel-Ziv Compression
So far we have accounted for the term $K(x|P^{*})$ only by very rough upper bounds rather than trying to approximate it. Here we describe a toy scenario in which we can detect anomalies using the Lempel-Ziv compression algorithm; more difficult scenarios may demand more powerful compression algorithms that are closer to general AI.
Consider the causal DAG: $X_{1}\to X_{2}\to\cdots\to X_{n},$ with the structural equations
$X_{1}=N_{1};\quad X_{j}=X_{j-1}+N_{j}\quad\text{for }j=2,\dots,n.$ (9)
Moreover, suppose that every $N_{j}$ for $j=1,\dots,n$ is drawn from the uniform distribution on the set of numbers in $[0,1]$ discretized to $d\gg 1$ digits of precision. By choosing the uniform distribution, the terms $-\log P(x_{j}\mid x_{j-1})$ become constant; moreover, we have $K(x_{j}\mid(x_{j-1},P)^{*})\stackrel{{\scriptstyle+}}{{=}}K(x_{j}\mid x_{j-1})$. Since the conditional distribution of each $X_{j}$, given its parent, is uniform over the (discretized) interval $[x_{j-1},\,x_{j-1}+1]$, the conditional randomness deficiency of $x_{j}$ reads
$\delta(x_{j}\mid x_{j-1})\stackrel{{\scriptstyle+}}{{=}}d\cdot\log 10-K(x_{j}\mid x_{j-1}).$
Suppose now we inject an anomaly at some node $j$ by setting $n_{j}$ to some numbers in $\{0,\dots,0.9\}$. As a result, $x_{j}$ and $x_{j-1}$ now coincide in at most $2$ digits, so that $K(x_{j}\mid x_{j-1})\stackrel{{\scriptstyle+}}{{=}}0$ and $\delta(x_{j}\mid x_{j-1})\stackrel{{\scriptstyle+}}{{=}}d\cdot\log 10$.
Following Steudel et al. [2010], we approximate $K(x_{j}\mid x_{j-1}^{*})\stackrel{{\scriptstyle+}}{{=}}K(x_{j},x_{j-1})-K(x_{j-1})$ by $R(x_{j},x_{j-1})-R(x_{j-1}),$ where $R$ denotes the length of a compressed encoding using the Lempel-Ziv algorithm. Whenever $n_{j}$ corrupts to a $1$-digit number, Lempel-Ziv recognizes that $x_{j}$ and $x_{j-1}$ coincide with respect to all but $1$ or $2$ digits. Thus, $R(x_{j},x_{j-1})\approx R(x_{j-1})$, which lets us infer the randomness deficiency of almost $d\log 10$.
We conducted experiments to verify our findings for $n=4$, with the noise uniformly drawn from numbers in $[0,1]$ with $d=10$ digits. We randomly chose one of the $4$ nodes as “root cause” and set all but one digit of its noise variables to zero. To detect the root cause, we selected the label $j$ that minimized Lempel-Ziv compression length and found the right one in $100$ out of $100$ runs.
From a theoretical point of view, the example shows that the joint observation can be anomalous (here in the sense of showing two variables whose digits coincide largely), although none of the variables show unexpected behaviour with respect to their marginal distribution. Hence, the root cause can only be found by inspecting which mechanism behaves unexpectedly.
7 Conclusion and Future Work
Algorithmic randomness deficiency offers a principled and flexible definition of outliers, without prior specification of the feature that exposes the anomaly. On a causal Bayesian network, we saw that the randomness deficiency of a joint observation decomposes along individual causal mechanisms, subject to the Independent Mechanisms Principle. This allows us to trace anomalous observations back to their root causes, identifying specific mechanisms that exhibit atypical behavior. Furthermore, we showed that weak outliers cannot be responsible for producing strong outliers, thus extending Levin’s law of randomness conservation. This foundational insight can help calibrating anomaly scores in a way that supports root cause analysis in complex systems.
Acknowledgments: DJ and YW would like to thank the authors of Li et al. [2024] for an interesting discussion about their Example 2.1.